OpenStudy (anonymous):

Given the rectangular equation xy=1, how do I write an equivalent polar equation?

OpenStudy (anonymous):

substitute: x=rcos(theta) y=rsin(theta)

OpenStudy (anonymous):

is that all I'm supposed to do? That isn't one of the answer choices

OpenStudy (anonymous):

you might have to do some algebra to get the answer choices...

OpenStudy (anonymous):

I guess this is lost on me I cannot figure out where to go.

OpenStudy (anonymous):

\[r^2\sin \theta \cos \theta=1\]\[\frac{r^2}{2}\sin (2\theta)=1\]

OpenStudy (anonymous):